OpenStudy (anonymous):

find the domain and sketch the graph.. f(x)=√x+9

OpenStudy (anonymous):

\[f(x)=\sqrt{x+9}?\]

OpenStudy (anonymous):

You cant take the square root of a negative number so, Whatever is in the square root has to be greater than or equal to 0 \[x+9\ge0\]

OpenStudy (anonymous):

yes, so the domain should be {x|x≥9} correct? how would that be graphed?

OpenStudy (anonymous):

The domain would be all numbers greater than or equal to -9. \[x \ge-9\] \[(-9,\infty)\] is your domain

OpenStudy (anonymous):

You know this is a transformation of the square root function, just take the graph of \[y=\sqrt{x}\] and translate it left 9 units

OpenStudy (anonymous):

okay i think i understand thank you !!!